OpenStudy (smurfy14):

Determine whether the function has a vertical asymptote or a removable decontinuity at x=-1. Determine whether the function has a vertical asymptote or a removable decontinuity at x=-1. @Mathematics

OpenStudy (anonymous):

find the limit of the function at the point where x = -1 :D

OpenStudy (smurfy14):

f(x)=(x^2-1)/(x+1)

OpenStudy (anonymous):

\[\lim_{x\rightarrow-1^-}{\frac{x^2-1}{x+1}} = \frac{0}{0^-} = 0\] \[\lim_{x\rightarrow-1^+}{\frac{x^2-1}{x+1}} = \frac{0}{0^+} = 0\]

OpenStudy (anonymous):

There seems to be a removable discontinuity at the point where x = -1

OpenStudy (smurfy14):

how do you figure that out?

OpenStudy (anonymous):

the left-hand and right-hand limits at x = -1 are the same number

OpenStudy (smurfy14):

oh ok thanks

OpenStudy (smurfy14):

what if it wasn the same?

OpenStudy (anonymous):

if it wasn't the same, then the discontinuity will not be removable.,

OpenStudy (anonymous):

it will be a V. asymptote :-D

OpenStudy (smurfy14):

oh ok! thanks again!

jimthompson5910 (jim_thompson5910):

another way is to simplify (x^2-1)/(x+1) to get (x^2-1)/(x+1) ((x+1)(x-1))/(x+1) x-1 So (x^2-1)/(x+1) = x-1 for all x such that x does not equal -1 So this means that there is a removable discontinuity at x = -1